40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Nov 30 11:44:57 2007, rob, Configuration, Electronics, ETMX oplev 
    Reply  Fri Nov 30 13:46:50 2007, rob, Configuration, Electronics, ETMX oplev dead again 
       Reply  Fri Nov 30 19:11:05 2007, rana, Configuration, Electronics, ETMX oplev dead again 
Message ID: 146     Entry time: Fri Nov 30 13:46:50 2007     In reply to: 145     Reply to this: 147
Author: rob 
Type: Configuration 
Category: Electronics 
Subject: ETMX oplev dead again 

Quote:
In the interests of getting the Xarm alignment script working again, I reset the local damping gains for the test masses to their previous known working values (1), then I noticed that the ETMX oplev was dead. Since the scripts use the oplev motion as a readback for the optic motion, this means the script was basically blindly swinging the optics around. Some monkeying around with swapping HeNe power supplies eventually led to the conclusion that the power strip is funky, since the laser works when plugged into another power strip. Even weirder, the HeNe and the power supply indicator light have some sort of XOR relationship going on. When one works, the other doesn't. Steve will sort out this confusion later; we're good for now.


Ech. The HeNe quit again. Let's replace it and see what happens.
ELOG V3.1.3-